OpenStudy (anonymous):

Use the law of logarithms to rewrite the expression below in a form with no logarithm of a product, quotient, or power.

OpenStudy (anonymous):



OpenStudy (anonymous):

I'll give medals :)

OpenStudy (whpalmer4):

\[\log_{6} \sqrt[7]{19}\]

OpenStudy (whpalmer4):

\[\log_{6} \sqrt[7]{19} =\frac{\ln (\sqrt[7]{19})}{\ln 6}=\frac{\frac{ 1 }{ 7 } \ln 19}{\ln 6} = \frac{\ln 19}{7 \ln 6}\]